OpenStudy (anonymous):

Choose the equation that could be used to find three consecutive integers whose sum is 36. n + (n + 2) + (n + 4) = 36 n + (n + 1) + (n + 3) = 36 n + (n + 1) + (n + 2) = 36 n + (n − 1) + (n − 3) = 36

zepdrix (zepdrix):

If we have an unknown number \(\large n\), what would be the next number after that? \(\large n+1\) Adding one would get us to the next number. The next number after that would be, \(\large n+2\) We simply added 1 again. This is how we would write 3 consecutive numbers. Do any of your options follow this idea?

OpenStudy (anonymous):

I am still confused, sorry :(

zepdrix (zepdrix):

So let's pretend that they wanted 3 consecutive numbers that add up to 6. Well that happens to be 1, 2 and 3, because, \(\large 1+(2)+(3)=6\) I put brackets around the numbers just to maybe help us understand what's going on. See our second number, the 2? There is another way we can write that. It's the first number plus 1. \[\large 1+(1+1)+(3)=6\]The 3 is also the first number, but plus 2.\[\large 1+(1+1)+(1+2)=6\] Now in our problem, our first number is not 1, it's instead an unknown number.\[\large n+(1+1)+(1+2)=\]But remember we were able to write all 3 numbers in terms of that first number.\[\large n+(n+1)+(n+2)=\] Does this kinda help with the idea of consecutive numbers? Or still confusing? :o

OpenStudy (anonymous):

No ;-;